Jonah Graham owns and operates The Green Thumb Company (GTC), which provides live plants and flower arrangements to professional offices. Jonah has fixed costs of $3,520 per month for office/greenhouse rent, advertising, and a delivery van. Variable costs for the plants, fertilizer, pots, and other supplies average $22 per job. GTC charges $62 per month for the average job.

Required:

1. How many jobs must GTC average each month to break even?

??? jobs per month

2. What is the operating income for GTC in a month with 85 jobs? Enter a net loss as a negative amount. $ What is the operating income for GTC in a month with 94 jobs?

$ ???

3. Jonah faces a tax rate equal to 20 percent. How many jobs must Jonah have per month to earn an after-tax income of $1,200? Round your answer to the nearest whole number of jobs.

??? jobs per month

4. Suppose that Jonah’s fixed costs increase to $3,652 per month and he decides to increase the price to $76 per job. What is the new break-even point in number of jobs per month? Round your answer to the nearest whole number of jobs.

??? jobs per month
Answer #1

1. Number of jobs must GTC average each month to break even is 88 jobs

Calculation:

Break even(in jobs) = Fixed cost per month ÷ Contribution per average job

= $3520/$40

= 88 Jobs

Contribution per average job = Sales per average job - variable cost per average job = 62-22 = $ 40

2.

85 Jobs 94 Jobs
Sales (No. of job x $62) 5270 5828
Less: Variable cost (No. of jobs x $22) -1870 -2068
Contribution 3400 3760
Less: Fixed cost -3520 -3520
Operating income/(loss) -120 240

3. Total number of jobs Jonah must have per month to earn an after-tax income of $1,200 are 126 jobs

calculation:

No. of jobs per month to earn an after tax income of $1,200 = (Profit before tax+fixed cost) ÷ Contribution per average job

= (1500+3520)/40

= 125.5 Jobs or 126 Jobs (Rounded)

Profit before tax = $1200/80% = $1500.

4. New break event point = New fixed cost÷ New contribution per average job

= 3,652/54

= 68 Jobs

New contribution per average job = $76-$22 = $54